OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $9 less than that for an adult's. The park offers half off all entry tickets during the off-peak season. The Sandlers paid a total of $132 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. : 132 = one-halfx + 2(x - 9) What is the regular price of a child's ticket? $60 $57 $51 $42

OpenStudy (anonymous):

@Night-Watcher

OpenStudy (texaschic101):

132 = 1/2x + 2 (x - 9) --- distribute through the parenthesis 132 = 1/2x + 2x - 18 -- add 18 to both sides 132 + 18 = 1/2x + 2x -- combine like terms 150 = 1/2x + 4/2x 150 = 5/2x 150/(5/2) = x 150 * 2/5 = x 300/5 = x 60 = x check.. 132 = 1/2(60) + 2(60 - 9) 132 = 30 + 2(51) 132 = 30 + 102 132 = 132 (correct) regular price of adults ticket (x) = $60 so if the regular price of a childs ticket is 9 less, then a childs ticket costs $51